OpenStudy (summersnow8):

The modulus of elasticity for a prosthetic material is 20 GPa. A 3 cm long sample of this material is circular in cross section with a radius of 1 cm. this sample is stretched 3.003 cm. what tensile force was applied to the material to create stretch? Supposedly the answer is 6,283 N

OpenStudy (michele_laino):

please you have to apply this formula: \[\frac{ F }{ S }=E \frac{ \Delta }{ L }\] where E is the Young modulus, Delta is your strtching, L is length of your sample, S is the cross section and F is the force applied

OpenStudy (summersnow8):

\[e = \sigma/\epsilon \] \[\sigma = F/A\] \[A = pir ^{2}\] = .000314 then I don't know what e equals in Pa

OpenStudy (summersnow8):

\[\epsilon = (.03003-.03)/(,03) = .00\]

OpenStudy (summersnow8):

.001*

OpenStudy (michele_laino):

I think Young modulus is equals to 20GPa*cross section

OpenStudy (michele_laino):

so E=6.3*10^3

OpenStudy (summersnow8):

huh? When i do this: \[20 GPa = F/ (.000314/.001) F= 6.28\]

OpenStudy (michele_laino):

sorry E = 6.3*10^6

OpenStudy (summersnow8):

I am lost

OpenStudy (michele_laino):

also me!

OpenStudy (summersnow8):

so i know A = .00314 & E= .001 then I have to plug it into the equation of \[\sigma=F/A\]

OpenStudy (michele_laino):

Please A=pi*10^(-4) m^2 do you agree?

OpenStudy (summersnow8):

Are you getting 6,283 as an answer?

OpenStudy (michele_laino):

no

OpenStudy (summersnow8):

that is suppose to be the answer to this problem, but I can't figure it out

OpenStudy (michele_laino):

Sorry, if i want the Young modulus I have to multiply this: 20*10^9*pi*10^(-4), do you agree?

OpenStudy (summersnow8):

NO, I don't understand where you are getting all that from

OpenStudy (michele_laino):

since your E have a dimension of a pressure

OpenStudy (michele_laino):

sorry E has a dimension...

OpenStudy (summersnow8):

can you just start from the beginning I am really confused. my prof said use the area of a circle... then solve for E , then plug those answers into the tensile force equation. I am getting tripped up on why 20 GPa changes... and how do we know

OpenStudy (michele_laino):

The formula for F, is: \[F=E \frac{ \Delta }{ L }A\] where E is the Young modulus, but E has the dimension of a force not of a pressure

OpenStudy (summersnow8):

okay, so how do you change GPa into the force of pressure then?

OpenStudy (michele_laino):

sorry E has the dimension of a pressure, sorry! I'm very sorry!

OpenStudy (michele_laino):

so I try to substitute your numerical data into the formula above

OpenStudy (michele_laino):

\[F=20*10^{9}\frac{ 3*10^{-2} }{ 3 }* \pi *10^{-4}=62.8*10^{3}N\]

OpenStudy (summersnow8):

uhhh, this makes no sense to me. Im so lost....

OpenStudy (michele_laino):

I know, nevertheless I have susbstitute your numerical data into formula above, and that formula is correct

OpenStudy (michele_laino):

1GPa=10^9 Pa, do you agree?

OpenStudy (summersnow8):

yes

OpenStudy (michele_laino):

stretching=3*10^(-3)cm, sorry not 3*10^(-2)cm, do you agree?

OpenStudy (michele_laino):

cross section A=3.14*10^(-4)m^2, do you agree?

OpenStudy (michele_laino):

since 1 cm=10^(-2) m

OpenStudy (summersnow8):

can you not do 10X 3 and put them in decimals. I'm getting confused

OpenStudy (summersnow8):

I know that A = pi (.01m) squared = .003 I know that E = (.03003 - .03)/(.03 = .001

OpenStudy (summersnow8):

so to find force= (20 GPa)/(.000314/.001) = 6.28

OpenStudy (summersnow8):

but i don't get how the answer is 6280

OpenStudy (michele_laino):

I got:

OpenStudy (michele_laino):

\[F=20*10^{9}\frac{ 3*10^{-5} }{ 3*10^{-2} }*\pi*10^{-4}=\] \[=62.8*10^{2}Newton=6,280Newton\]

OpenStudy (michele_laino):

please keep in mind that youe stretching is 3.003-3=0.003 cm

OpenStudy (summersnow8):

yeah i know but i don't get your method. i don't get the 20 to the 9th and 3 times 10 to the -5th and why you are multiplying by pi and 10 to the -4

OpenStudy (michele_laino):

because cross section is: \[\pi*1cm ^{2}=\pi*10^{-4}m ^{2}\] since: \[1m ^{2}=10^{4}cm ^{2}\]

OpenStudy (summersnow8):

i still don't understand. I'm just going to give up. i said use decimals and you keep using 10 to the 4th and stuff. just forget it

OpenStudy (michele_laino):

cross section A = pi*1cm^2 do you agree?

OpenStudy (summersnow8):

yes

OpenStudy (michele_laino):

please keep in mind you have exprees all your quantity in MKS unit

OpenStudy (michele_laino):

now 1cm=0.01m do you agree?

OpenStudy (summersnow8):

yea

OpenStudy (michele_laino):

and 1 cm^2=0.0001 m^2 =10^(-4) m^2 do you agree?

OpenStudy (summersnow8):

yeah, but i prefer decimals

OpenStudy (michele_laino):

ok! then please use this: 1 cm^2=0.0001 m^2 and stretching =0.00003 m length=0.03 m

OpenStudy (summersnow8):

so yours is really saying: F=20000000000 (.0003/.03) x pi .001 ?

OpenStudy (summersnow8):

i don't know why you multiply the answer by 10^2 newtons

OpenStudy (michele_laino):

no, please note that A =pi*0.0001 m^2 stretching =0.00003 m

OpenStudy (michele_laino):

because I use the so called scientific notation, pratically the notation in which appear power of 10

OpenStudy (michele_laino):

oops sorry appear powers of 10...

OpenStudy (michele_laino):

did you get the right answer?

OpenStudy (summersnow8):

? E= (.03003 - .03) / (.03) = .001 A= pi (.01)^2 = .000314

OpenStudy (michele_laino):

ok!

OpenStudy (summersnow8):

no

OpenStudy (michele_laino):

ok!

OpenStudy (summersnow8):

i don't get it

OpenStudy (summersnow8):

what if the problem had 20 pa instead of GPa what would the answer be then

OpenStudy (michele_laino):

using decimal I can write: \[F=20000000000*\frac{ 0.00003 }{ 0.03 }*\pi*0.0001\]

OpenStudy (summersnow8):

how did you get .00003

OpenStudy (michele_laino):

because your stretching is equals to final length-initial length, namely final length=3.003 cm and initial length=3 cm, so stretching=3.003-3=0.003 cm, finally stretching = 0.00003 meters

OpenStudy (michele_laino):

@Summersnow8

OpenStudy (summersnow8):

ok

OpenStudy (michele_laino):

did yo get your answer, using decimal and not power of 10 instead?

OpenStudy (summersnow8):

no, i don't get why we don't just use 20....

OpenStudy (summersnow8):

im stuck on that

OpenStudy (michele_laino):

It is very simple, now I rewrite your answer using only decimals:

OpenStudy (michele_laino):

\[F=20000000000*\frac{ 0.00003 }{ 0.03 }*\pi*0.0001\]

OpenStudy (michele_laino):

where all quantities are measured with MKS units, namely length in meters, section in squared meters Young modulus in Pascal force in Newton

OpenStudy (michele_laino):

please complete the calculus above

OpenStudy (summersnow8):

yeah I know. but how do i get the answer to Pa

OpenStudy (summersnow8):

and what if 20 is already in Pa do I have to convert?

OpenStudy (summersnow8):

i got 3283.18

OpenStudy (michele_laino):

please the answer is a force not a pressure

OpenStudy (summersnow8):

6283

OpenStudy (michele_laino):

did you get 6283?

OpenStudy (summersnow8):

yeah

OpenStudy (michele_laino):

ok! perfect! well done!

OpenStudy (summersnow8):

that was too hard

OpenStudy (summersnow8):

so if 20 Pa was in place of 20GPa i would just use 20?

OpenStudy (michele_laino):

that's right! because Pa is an unit of measure that belongs to MKS system of measure units

OpenStudy (summersnow8):

so then the answer would be 6.28

OpenStudy (summersnow8):

@Michele_Laino

OpenStudy (michele_laino):

no, sorry you will get a force which is 0.000000001 times the force you got now!

OpenStudy (summersnow8):

huh? so......... it would be .000000002

OpenStudy (summersnow8):

.00000000628?

OpenStudy (michele_laino):

no, sorry 0.000000001*6280= 0.00000628 Newton

OpenStudy (summersnow8):

how do you know how many decimals?

OpenStudy (michele_laino):

because if the Young modulus is 20Pa then it is 10^-9 times the preceding Young modulus which was 10GPa, so our answer will be 10^-9 times preceding answer

OpenStudy (michele_laino):

sorry ...whic was 20 GPa, ...

OpenStudy (summersnow8):

I get when we use 20 GPa.... but lets say for this problem instead of 20 GPa it says 20 Pa.... what would the answer be then

OpenStudy (michele_laino):

as I said before, your answer will be: F=0.00000628 Newton

OpenStudy (summersnow8):

whats the conversion then? instead of using 20000000000